Keller Construction is considering two new investments. Project E calls for the purchase of earthmoving equipment. Project H represents an investment in a hydraulic lift. Keller wishes to use a net present value profile in comparing the projects. The investment and cash flow patterns are as follows: Use Appendix B for an approximate answer but calculate your final answer using the formula and financial calculator methods.

Project E Project H
($37,000 Investment) ($35,000 Investment)
Year Cash Flow Year Cash Flow
1 $ 9,000 1 $ 17,000
2 12,000 2 18,000
3 18,000 3 17,000
4 20,000

a. Determine the net present value of the projects based on a zero percent discount rate.

NPV
Project E
Project H

b. Determine the net present value of the projects based on a discount rate of 9 percent. (Do not round intermediate calculations and round your answers to 2 decimal places.)

NPV
Project E
Project H

c. If the projects are not mutually exclusive, which project(s) would you accept if the discount rate is 9 percent?
Answer #1

A B H a) year 0 C D E G discount Project E rate 0% value Project H value - 37000 1 - 37000 - 35000 35000 9000 1 9000 17000 17

discount Project E rate 0% -37000 1 9000 =+D2/1 12000 =+D3/1 18000 =+D4/1 20000 =+D5/1 value =+C2D2 =+C3D3 =+C4D4 =+C56D5 =+C

PLEase ask if any query.